P6.3A Vista Company Inc. had a beginning inventory of 100 units of Product RST at a cost of $8 per unit. During the year, purchases were:

Feb. 20    600 units at $ 9    Aug. 12    400 units at $11
May  5 500 units at $10 Dec.  8 100 units at $12

Vista Company uses a periodic inventory system. Sales totaled 1,500 units.

Instructions

a. Determine the cost of goods available for sale.

b. Determine the ending inventory and the cost of goods sold under each of the assumed cost flow methods (FIFO, LIFO, and average-cost). Prove the accuracy of the cost of goods sold under the FIFO and LIFO methods. (Round average unit cost to three decimal places.)
Answer #1
Unit Unit cost Total cost
Beginning inventory 100 8 800
Purchase 600 9 5400
Purchase 2 500 10 5000
Purchase 3 400 11 4400
Purchase 4 100 12 1200
Total 1700 16800

a) Cost of goods available for sale = $16800

b) Calculate following

FIFO LIFO Average cost
Ending inventory (100*12+100*11) = 2300 (800+900) = 1700 16800/1700 = 9.882*200 = 1976
Cost of goods sold 16800-2300 = 14500 16800-1700 = 15100 16800-1976 = 14824